### v2

# RIDE: Difficulty Evolving Perturbation with Item Response Theory for Mathematical Reasoning

Abstract
Large language models (LLMs) achieve high performance on mathematical reasoning, but these can be inflated by training data leakage or superficial pattern matching rather than genuine reasoning. To this end, an adversarial perturbation–based evaluation is needed to measure true mathematical reasoning ability. Current rule-based perturbation methods often generate ill-posed questions and impede the systematic evaluation of question difficulty and the evolution of benchmarks. To bridge this gap, we propose RIDE, a novel adversarial question-rewriting framework that leverages Item Response Theory (IRT) to rigorously measure question difficulty and to generate intrinsically more challenging, well-posed variations of mathematical questions. We employ $35$ LLMs to simulate students and build a difficulty ranker from their responses. This ranker provides a reward signal during reinforcement learning and guides a question-rewriting model to reformulate existing questions across difficulty levels. Applying RIDE to competition-level mathematical benchmarks yields perturbed versions that degrade advanced LLM performance, with experiments showing an average $21.73\%$ drop across $26$ models, thereby exposing limited robustness in mathematical reasoning and confirming the validity of our evaluation approach.
Machine Learning, ICML
1 Introduction
Mathematical reasoning is regarded as a crucial testament for evaluating the reasoning capabilities of large language models (LLMs). Recent works like DeepSeek-Math (Shao et al., 2024) and Qwen2.5-Math (Yang et al., 2024) clearly demonstrate the benefits of reinforcement learning for mathematical reasoning. Nevertheless, strong performance in mathematical reasoning may not directly reflect the model’s reasoning ability (Wu et al., 2025). It may instead result from data contamination during pretraining (Golchin and Surdeanu, 2024) that leads to memorization-based answers (Zhang et al., 2024; Li et al., 2025), or from the model relying primarily on superficial pattern matching without understanding mathematical knowledge (Li et al., 2024; Mirzadeh et al., 2025). This phenomenon undermines the reliability of LLMs as genuine mathematical reasoners.
To address this issue, there is an urgent need for a more adversarial evaluation benchmark that can faithfully characterize the robustness of mathematical reasoning. For example, a rule-based rewriting method is proposed (Li et al., 2024; Zhou et al., 2024), which applies operations such as numerical substitution and symbol modification to reformulate questions. This method has been shown to reduce model performance. However, rule-based approaches may render the rewritten questions ill-posed and unsolvable (Xue et al., 2025), thereby undermining their validity as reliable benchmarks. Moreover, previous work has focused mainly on low-difficulty datasets (Cobbe et al., 2021), with limited attempts at more challenging data. Based on this, we identify the following challenges: (1) Robust Adversarial Perturbation. Robust perturbations to the data need to both degrade the model performance and remain credible and reasonable, which requires shifting from injecting noise to reformulating questions into formats that are inherently more challenging for the model to resolve correctly. Moreover, the perturbation should be unpredictable for changing, and simple rules cannot capture the evolution. (2) More Challenging Data. The perturbation of data should not be limited to the same difficulty level of mathematical reasoning. It is also necessary to explore its effectiveness on difficulty evolution, ensuring reasonable perturbations can be applied to mathematical competition questions. These challenges require us to perturb existing benchmarks in ways that substantively increase the intrinsic difficulty of the questions (Huang et al., 2025). A question’s difficulty is often defined by whether a model answers correctly (Lan et al., 2024), but this definition depends heavily on the selected models, and majority voting across these models ignores the individual differences in capabilities. Besides, defining difficulty by the steps of the reasoning chain (Yu et al., 2025) is easily misled by shortcut reasoning and does not account for the numerical computation complexity specific to mathematical reasoning.
Figure 1: Difficulty evolution via RIDE. The original question (left) is successfully answered, whereas the rewritten question (right) with added constraints triggers a reasoning error of the Qwen3-30B-A3B model.
To this end, we propose RIDE, a benchmark perturbation rewriting framework that incorporates Item Response Theory (IRT) to measure question difficulty (Vania et al., 2021; Scarlatos et al., 2025). Specifically, $35$ cutting-edge LLMs are treated as students to answer $2,000$ questions sampled from the high-difficulty mathematical dataset DeepMath (He et al., 2025). A statistical model jointly characterizes the latent abilities of the student models and the difficulty parameters of the questions. This difficulty is not human cognitive complexity but a model-centric metric representing the inherent challenge a specific question format poses to the model’s ability to respond correctly. Based on the IRT-estimated difficulty, a pairwise ranker is trained: given question embedding vectors, it outputs the probability that one question is more difficult than another. The ranker’s output further serves as a difficulty reward signal for reinforcement learning on Qwen3-8B, resulting in the question rewriting model RIDE-8B. Applying reinforcement learning to difficulty-evolving question rewriting encourages the model to adaptively generate questions with higher difficulty, thus avoiding superficial modifications, while further enhancing consistency between the rewritten questions and their answers by setting a correctness reward. RIDE-8B is then applied to rewrite competition-level mathematical benchmarks such as AIME-24 and AMC-23, achieving difficulty evolution for these datasets. Experiments demonstrate that the perturbed benchmarks induce performance drops in frontier reasoning models, including GPT-5, DeepSeek-V3.1, and others, thereby enabling robust evaluation and comparison of advanced models’ mathematical reasoning abilities (as shown in Figure 1). In addition, RIDE can be used as a data augmentation method to enhance training data and improve mathematical reasoning performance. Our contributions are as follows:
  • We propose a benchmark perturbation framework with IRT to evaluate robustness in mathematical reasoning. We measure robustness via performance degradation and observe that most LLMs are sensitive to perturbations, indicating poor robustness.
  • We propose an IRT-based difficulty ranker and introduce a reinforcement learning scheme for training a question-rewriting model. We also release our question rewriting model (RIDE-8B), two rewritten competition-level benchmarks (RIDE-AIME and RIDE-AMC), and a training dataset of over 10K rewritten questions (RIDE-DeepMath).
2 Related Work
Robustness Evaluation. Existing LLM robustness evaluations (Ul Abedin et al., 2025) often use controlled perturbations to test whether models truly understand tasks rather than memorize surface patterns (Li et al., 2024; Yang et al., 2025a). A common practice is to construct original items and perturbed counterparts around the same semantics and compare the resulting performance drop (Goel et al., 2021; Chang et al., 2023; Hou et al., 2025; Yang et al., 2025b; Sun et al., 2025). For example, GSM-PLUS (Li et al., 2024) applies rule-based perturbations such as numerical substitution (Shrestha et al., 2025) and operation modification, revealing substantial degradation in many LLMs. Unlike rule-based methods, we adapt a reinforcement learning-based rewriting approach to generate perturbations, producing better-formed data, avoiding unsolvable or multiple-solution cases, and increasing unpredictability.
Item Response Theory (IRT). Originating from educational measurement, IRT is widely applied in AI for automated item difficulty estimation (Feng et al., 2025; Chen and Shiu, 2025) and efficient evaluation (Polo et al., 2024; Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022; Xu et al., 2025b). IRT depends on real student responses, which are often difficult to obtain due to privacy constraints. Prior work addresses this by simulating response matrices (Lalor et al., 2019; Truong et al., 2025) or predicting difficulty from text (Martínez-Plumed_Castellano_Monserrat-Aranda_Hernández-Orallo_2022), which leads to high computational costs (Truong et al., 2025). We mitigate these issues via data augmentation to reduce the number of required student models and reframing difficulty estimation as a pairwise ranking task to reduce numerical prediction errors.
3 Preliminary
We briefly introduce the IRT-based difficulty in our method. Consider a test taker with fixed ability $\theta$ and a math reasoning question $q$ with a difficulty score $d$. The response $y$ is modeled as a Bernoulli random variable that indicates correctness, where $y=1$ denotes a correct answer and $y=0$ denotes an incorrect answer. We introduce the Rasch model (Rasch, 1993), a foundational one-parameter (1-PL) IRT model in which the probability of a correct response is the logistic function of the difference between the test taker’s ability $\theta$ and the item’s difficulty $d$:
$\Pr(y=1\mid\theta,d)=\sigma(\theta-d)=\frac{1}{1+e^{-(\theta-d)}}$
Given a binary response matrix $\mathbf{Y}\in\{0,1\}^{M\times N}$, where $M$ and $N$ denote the number of test takers and questions, respectively. Each entry $Y_{ij}$ indicates the $i$-th test taker’s answer to the $j$-th question. Based on this response matrix, one can estimate the ability parameters $\theta$ and the difficulty parameters $d$ using a variety of statistical inference methods, such as the Bayesian inference via Markov chain Monte Carlo (MCMC) sampling (Chen et al., 2021) and variational inference (VI) (Blei et al., 2017). In this paper, we apply the VI method provided by py-irt (Lalor and Rodriguez, 2023). Concretely, we fit the Rasch model via stochastic variational inference on the response matrix with a mean-field Normal guide for $\theta$ and $d$, maximizing the evidence lower bound $\mathcal{L}$:
$\mathrm{KL}(g\|p)=\sum_{i=1}^{M}\mathrm{KL}\!\big(g(\theta_{i})\,\|\,p(\theta_{i})\big)+\sum_{j=1}^{N}\mathrm{KL}\!\big(g(d_{j})\,\|\,p(d_{j})\big)$
$\max_{g}\ \mathcal{L}=$ $\sum_{i=1}^{M}\sum_{j=1}^{N}\mathbb{E}_{g}\!\left[Y_{ij}(\theta_{i}-d_{j})-\log\!\big(1+e^{\theta_{i}-d_{j}}\big)\right]$
$-\mathrm{KL}\!\big(g\,\|\,p\big)$
Here, $g$ denotes the variational joint distribution over the latent parameters and $p$ denotes the prior joint distribution over $(\theta,d)$. $\mathrm{KL}(g\|p)$ denotes the Kullback-Leibler divergence from $g$ to $p$. We take posterior means as point estimates from $g$ to obtain IRT-based difficulty $d$ for each question.
4 Method
In this section, we elaborate on the details of RIDE-Math. Inspired by the IRT-based difficulty estimation method with LLMs acting as students (Truong et al., 2025), we construct a difficulty ranker for math questions by performing parameter estimation described in Section 3. This ranker serves as a reward model by providing difficulty reward signals, which are used to guide reinforcement learning on the existing high-difficulty dataset DeepMath (He et al., 2025). Through this rewriting process, we are able to generate difficulty in evolving math reasoning data. The overall architecture of our method is shown in Figure 2.
Figure 2: The overall architecture of RIDE. RIDE trains a difficulty ranker based on IRT estimations to provide reward signals, guiding reinforcement learning to drive the LLM to rewrite original math questions into variations that expose model’s reasoning vulnerabilities.
4.1 IRT-guided Difficulty Ranker
4.1.1 IRT Parameter Estimation
We draw $N=2,000$ math questions from the DeepMath dataset to construct a quiz. We then employ $M=35$ cutting-edge LLMs with different parameter scales and architectures as students to solve these questions, thereby constructing a response matrix $\mathbf{Y}\in\{0,1\}^{M\times N}$. Specifically, for a math question $j$, the response vector can be described as
$\mathbf{v}^{j}=[\mathbb{I}[g^{j}=y_{1}],\mathbb{I}[g^{j}=y_{2}]...,\mathbb{I}[g^{j}=y_{M}]]^{\intercal}$
Here, $\mathbb{I}$ denotes the 010–1 indicator function, $g^{j}$ denotes the ground-truth answer to the $j$-th question, $y_{i}$ represents the response of the $i$-th student model. After obtaining the response vectors for all $N$ questions, we combine them to form the response matrix $Y$:
$\mathbf{Y}=\bigl[\,\mathbf{v}^{1},\;\mathbf{v}^{2},\;\ldots,\;\mathbf{v}^{N}\,\bigr]$
We perform IRT estimation based on the response matrix $\mathbf{Y}$. Following the parameter estimation method described in Section 3, we obtain an IRT-estimated difficulty $d$ for each question and construct $N$ question-difficulty pairs $(q,d)$ for training the difficulty ranker, where $q$ denotes the question.
4.1.2 Pairwise Difficulty Ranker
In traditional IRT, estimating the difficulty of a new question requires collecting fresh student response data. To endow IRT difficulty estimation with generalization to unseen questions, we adopt an amortized approach (Truong et al., 2025), which extracts content features via an embedding model and learn a mapping from question content to difficulty value. However, due to the domain-specific symbolic semantics of math questions (Mirzadeh et al., 2025), directly regressing from embeddings to a scalar difficulty may result in unreliable performance. Therefore, instead of regressing the difficulty value, we frame the task as a pairwise ranking question, training a ranker to identify the more difficult question from a pair of question embeddings. We use OpenAI’s text-embedding-3-large  [^1] https://platform.openai.com/docs/models/text-embedding-3-large model to obtain embedding vectors for each question-difficulty pair $(q,d)$, creating $N$ question-embedding-difficulty triplets as $(q,e,d)$. We then group these pairs by fine-grained topics (Gao et al., 2025) like algebra, geometry, and so on. Within each subcategory, ranking data are generated by forming pairwise combinations of questions. Specifically, for any two questions $(q_{i},q_{j})$ within the same subcategory, we create the ordered pair $(e_{i},e_{j},r_{ij})$ and its symmetric counterpart $(e_{j},e_{i},1-r_{ij})$ to reduce positional bias, where $r_{ij}$ denotes to the ranking label: it equals to $1$ if the question $i$’s difficulty $d_{i}$ exceeds the question $j$’s difficulty $d_{j}$, and 0 otherwise. To account for the influence of the magnitude gap between difficulty scores $d$, we assign a weight to each ranking pair. For $(e_{i},e_{j},r_{ij})$, the weight is $w_{ij}=\sqrt{\lvert d_{i}-d_{j}\rvert}$. We use these pairwise samples to train a multilayer perceptron (MLP) as the difficulty ranker. The input $\mathbf{x}$ of MLP is a fusion of absolute and relative features, which is concatenated as:
[…]
Weighted binary cross-entropy loss is used as the optimization target to train an MLP with hyperparameter $\phi$. The MLP ranker outputs the logit and then calculated $p(i,j)$, which measures the probability that question $i$ is more difficult than question $j$:
$p(i,j)=\sigma(\operatorname{MLP}_{\bm{\phi}}(\mathbf{x}_{ij}))$
$\mathcal{L}_{\mathrm{BCE}}=$ $\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\mathcal{B}}w_{ij}\Big(-r_{ij}\log p(i,j)$
$-(1-r_{ij})\log\!\big(1-p(i,j)\big)\Big)$
$p_{ij}$ will be used to form the difficulty reward for rewriting questions in reinforcement learning.
4.2 Reinforcement Learning for Difficulty Evolution
We use Qwen3-8B (Team, 2025) as the base model and train a math question rewriting model via Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). The SFT stage primarily serves as a cold start, which equips the base model with an initial ability to rewrite original math questions $q$ into more difficult ones $q^{\prime}$ and standardizes the output format. We curate $2,000$ high-quality rewritten questions $\mathcal{D}$ by distilling and filtering GPT-5-mini  [^2] https://openai.com/ outputs, and used this corpus to perform full-parameter fine-tuning of Qwen3-8B:
$\theta_{\text{SFT}}\xleftarrow{}\underset{\theta}{\mathrm{argmin}}\;\mathbb{E}_{(q,q^{\prime})\in\mathcal{D}}\big[-\mathrm{log}(\mathcal{P}_{\theta}(q^{\prime}\mid q))\big]$
$\theta$ and $\theta_{SFT}$ denote to the original parameter of the base model and the new parameter after full-parameter fine-tuning, respectively.
In the RL stage, we use the Group Sequence Policy Optimization (GSPO) algorithm (Zheng et al., 2025) for optimization. Starting from SFT model with parameter $\theta_{\text{SFT}}$, for each $q$ we sample a group $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}\sim\mathcal{P}_{\theta}(\cdot|q)$ with group size $K$, compute rewards $R(q,q_{k}^{\prime})$ and group-normalized advantages, and update model parameter $\theta$ with GSPO. Our rewards consist of:
  • Difficulty Reward. The difficulty reward measures whether the rewritten question is more difficult than the original. We obtain embedding vectors $e$ and $e^{\prime}$ for the original question $q$ and its rewrite $q^{\prime}$, respectively, and concatenate them as the input to the difficulty ranker. The ranker returns $p(q^{\prime},q)$ and its reverse $p(q,q^{\prime})$, which respectively describe the probabilities that $q^{\prime}$ is harder than $q$ and vice versa. The difficulty reward is defined as:
    $R_{\text{diff}}\;=\;p(q^{\prime},q)\;-\;p(q,q^{\prime})\,$
  • Correctness Reward. Due to the base model’s limited reasoning ability, the rewritten questions may contain errors. We use a teacher model GPT-5-mini to verify each rewrite and output a correctness reward $R_{\text{cor}}\in\{-1,1\}$, with 1-1 for incorrect and $1$ for correct.
Given reward weights $\alpha$ and $\beta$, the final reward $R$ is defined as their weighted sum. For each $q$ with samples $\mathcal{G}(q)=\{q_{k}^{\prime}\}_{k=1}^{K}$, let $R_{k}\equiv R(q,q_{k}^{\prime})$ and compute advantage $A_{k}$:
$R=\alpha R_{\text{diff}}+\beta R_{\text{cor}},\quad A_{k}=\frac{R_{k}-\tfrac{1}{K}\sum_{k=1}^{K}R_{k}}{\mathrm{std}(R_{1:K})},$
where $\operatorname{std}(R_{1:K})$ denotes the standard deviation over the group $\{R_{i}\}_{i=1}^{K}$. Let $\theta_{\text{old}}$ be the policy for sampling $\mathcal{G}(q)$, initialized as $\theta_{\text{SFT}}$ and updated online to $\theta$. The sequence-level importance ratio:
${\rho}_{k}(\theta)$ $=\left(\frac{\mathcal{P}_{\theta}(q_{k}^{\prime}\mid q)}{\mathcal{P}_{\theta_{\text{old}}}(q_{k}^{\prime}\mid q)}\right)^{\!1/|q_{k}^{\prime}|}$
$=\exp\!\left(\frac{1}{|q_{k}^{\prime}|}\sum_{t=1}^{|q_{k}^{\prime}|}\log\frac{\pi_{\theta}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}{\pi_{\theta_{\text{old}}}\!\big(q^{\prime}_{k,t}\mid q,q^{\prime}_{k,<t}\big)}\right)$
Here, $t$ indexes positions in the generated response $q_{k}^{\prime}$; $q^{\prime}_{k,t}$ denotes the $t$-th response token and $q^{\prime}_{k,<t}$ is its prefix. The normalization length $|q_{k}^{\prime}|$ counts response tokens only, and accordingly $\log\mathcal{P}_{\theta}(q_{k}^{\prime}\!\mid q)$ is computed as the sum of token-level log probabilities over those response positions. We maximize the clipped surrogate:
$\mathcal{L}_{\mathrm{GSPO}}(\theta)=$ $\mathbb{E}_{q\in\mathcal{D}}\Bigg[\frac{1}{K}\sum_{k=1}^{K}\min\Big(\rho_{k}(\theta)A_{k},$
$\text{clip}\big(\rho_{k}(\theta),1-\varepsilon,1+\varepsilon\big)A_{k}\Big)\Bigg]$
At iteration $i$, we set $\theta_{\text{old}}=\theta^{(i)}$; for each source question $q$, we sample a group $\mathcal{G}(q)$ and compute the component rewards $R_{\mathrm{diff}},\,R_{\mathrm{cor}},\,R_{\mathrm{key}}$ and their mixture $R$, form the group-normalized advantages $A_{k}$, and then update the parameters by maximizing the GSPO surrogate:
$\theta^{(i+1)}=\arg\max_{\theta}\mathcal{L}_{\mathrm{GSPO}}(\theta)$
Once the reward converges, we obtain difficulty-aware math rewriting model RIDE-8B. Considering that existing mathematical datasets may not provide complete solutions, we only take the original question $q_{\text{original}}$ and its answer $a_{\text{original}}$ from benchmarks or training data as input. The corresponding prompt is shown in Appendix A.3. Instead of prescribing specific rules, the prompt provides high-level guidance for increasing difficulty, allowing the model to freely rewrite questions, thereby creating unpredictable paths of difficulty progression. Subsequently, we take the prompt context as input into RIDE-8B, which infers the rewritten question $q_{\text{rewrite}}$, solution $s_{\text{rewrite}}$, and answer $a_{\text{rewrite}}$:
$(q_{\text{original}},a_{\text{original}})\;\;\xrightarrow{\;\;\text{RIDE-8B}\;\;}(q_{\text{rewrite}},s_{\text{rewrite}},a_{\text{rewrite}}).$
Including the solution $s_{\text{rewrite}}$ as a part of rewriting output not only strengthens the consistency between the rewritten question and its answer but also serves as a chain-of-thought signal, which can facilitate further training of models to improve their mathematical reasoning ability.
5 Experiment
5.1 Experiment Setup
We use our rewriting model RIDE-8B to perform multi-round difficulty evolution on the competition-level mathematical-reasoning benchmarks AMC-23 and AIME-24, which contain $40$ and $30$ questions, respectively. Using GPT-5, we filter out questions that are unsolvable or have incorrect answers to construct the final perturbed benchmarks RIDE-AMC and RIDE-AIME, which are further manually checked to ensure answer accuracy, with annotators holding at least a bachelor’s degree. We evaluate the pass@1 performance of cutting-edge open-source and proprietary models in the zero-shot setting. In addition, we extract data from the DeepMath dataset for rewriting, and after filtering with DeepSeek-V3.1 (DeepSeek-AI, 2024), we obtain 1010K high-quality mathematical reasoning samples, referred to as RIDE-DeepMath, which serve as training data to improve the models’ ability for mathematical reasoning.
Model Params AIME-24 RIDE-AIME PDR (↓%) AMC-23 RIDE-AMC PDR (↓%)
Proprietary Models
Claude-4-Sonnet - 40.00 36.67 8.32 85.00 82.50 2.94
Claude-4.1-Opus - 53.33 36.67 31.24 87.50 82.50 5.71
Gemini-2.0-Flash - 30.00 30.00 0.00 85.00 80.00 5.88
Gemini-2.5-Pro - 90.00 90.00 0.00 100.00 100.00 0.00
GPT-3.5-turbo - 10.00 3.33 66.70 55.00 50.00 9.09
GPT-o3-mini - 76.67 76.67 0.00 100.00 97.50 2.50
GPT-5 - 93.33 86.67 7.14 100.00 100.00 0.00
Grok-4 - 100.00 96.67 3.33 100.00 95.00 5.00
Open-source Models
DeepSeek-V3.1 671B 60.00 56.67 5.55 87.50 85.00 2.86
DeepSeek-V3.1-Thinking 671B 93.33 83.33 10.71 100.00 100.00 0.00
GLM-4.5 355B 93.33 80.00 14.28 100.00 92.50 7.50
Kimi-K2-Instruct 1TB 73.33 56.67 22.72 90.00 85.00 5.56
Llama2-70B 70B 26.67 3.33 87.51 40.00 30.00 25.00
Llama3.1-70B 70B 20.00 6.67 66.65 47.50 30.00 36.84
Llama3.1-405B 405B 23.33 6.67 71.41 47.50 35.00 26.32
Llama4-Scout 109B 30.00 3.33 88.90 57.50 45.00 21.74
Mistral-large 123B 30.00 20.00 33.33 70.00 60.00 14.29
Qwen2.5-7B 7B 10.00 10.00 0.00 50.00 30.00 40.00
Qwen2.5-Math-7B 7B 16.67 3.33 80.02 60.00 32.50 45.83
Qwen2.5-Math-72B 72B 26.67 10.00 62.50 70.00 60.00 14.29
Qwen2.5-72B 72B 16.67 10.00 40.01 70.00 52.50 25.00
Qwen3-0.6B 0.6B 10.00 3.33 66.70 37.50 20.00 46.67
Qwen3-30B-A3B 30B 83.33 80.00 4.00 80.00 72.50 9.38
Table 1: Comparison of the model’s pass@1 performance before and after benchmark perturbation. After benchmark perturbation, most models show a significant performance drop. The top-33 models with the largest average performance decline are highlighted in blue.
5.2 Main Results
We evaluate $23$ LLMs comprising $8$ proprietary models (including GPT-5 and Grok-4) and $15$ open-source models (including DeepSeek-V3.1 and Kimi-K2-Instruct (Team et al., 2025)), spanning parameter counts from 0.60.6B to 11TB and covering multiple model series. We report pass@1 performance on the original AMC-23 and AIME-24 benchmarks, and then apply the same prompts to our perturbed benchmarks, RIDE-AMC and RIDE-AIME, again recording pass@1 performance. To quantify robustness, we compute the performance drop rate (PDR) (Li et al., 2024):
$\mathrm{PDR}\;=\;1\;-\;\frac{\tfrac{1}{\lvert\mathcal{D}_{a}\rvert}\sum_{(x,y)\in\mathcal{D}_{a}}\mathbb{I}\!\left[\mathrm{LM}(x)=y\right]}{\tfrac{1}{\lvert\mathcal{D}\rvert}\sum_{(x,y)\in\mathcal{D}}\mathbb{I}\left[\mathrm{LM}(x)=y\right]},$
where $x$ and $y$ denote the question and the corresponding answer, respectively, $\mathcal{D}$ is the original benchmark (AMC-23 or AIME-24), $\mathcal{D}_{a}$ is the corresponding perturbed benchmark (RIDE-AMC or RIDE-AIME), $\mathrm{LM}(\cdot)$ is the model’s prediction, and $\mathbb{I}[\cdot]$ is the indicator function. Lower PDR indicates greater robustness. Results are presented in Table 1.
For AMC-23 and RIDE-AMC we use Qwen3’s non-thinking mode, whereas for AIME-24 and RIDE-AIME we use Qwen3’s thinking mode. Results show that our difficulty evolution perturbation method leads to performance drops in the vast majority of both proprietary and open-source models, even state-of-the-art models including GPT-5 and DeepSeek-V3.1-Thinking exhibit declines, which shows the validation of our method. On the more challenging benchmark AIME-24, the average PDR is higher than on AMC-23. Across all models, the mean PDR reached 21.7321.73%. From the model perspective, proprietary models generally have lower PDRs, indicating better robustness. Open-source models showed substantial variation, and several LLaMA and Qwen series models suffered large declines, with the highest PDR reaching 88.9088.90%.
6 Analysis
In this subsection, we conduct more experiments to explore and analyze the effectiveness and impact of our method.
Pass@n Performance. To verify the stability of our method against model performance perturbations, we evaluate Qwen2.5-Math-72B and Llama-4-Scout on AIME-24, AMC-23, and their perturbed RIDE-AIME and RIDE-AMC using pass@n (n = 1, 2, 4, 8). Results (Figure 3) show that accuracy improves as n increases, but performance on perturbed benchmarks remains notably lower, confirming the stability of our method in testing robustness. Moreover, Qwen2.5-Math-72B exhibits greater stability, with a relatively small gap between its pass@1 and pass@8 scores.
Figure 3: Comparison of pass@n performance before and after benchmark perturbation. Despite scaling improvements with $n$, the persistent performance gap confirms RIDE’s stability.
Win Rate Comparison of RIDE and Other methods. To evaluate rewrite quality, we design prompts assessing completeness, clarity, conceptual consistency, structural perturbations, and avoidance of shortcuts (Appendix A.3). GPT-5 acts as an evaluator to compare the outputs from different methods. Each pair is judged twice with swapped order to mitigate position bias. We compute two win rate metrics: (1) Average win rate—the proportion of times a method is selected across both evaluations; (2) Consistent win rate—the proportion of questions where a method is selected in both evaluations. Ties are allowed in the evaluator’s output. We run three rounds and report average results. We compare our method with the following method: (1) Rule-based method. Following GSM-Plus (Li et al., 2024), we use Qwen3-8B to generate perturbed AMC-23 and AIME-24 data with the same prompting strategy as rule-based baseline. (2) Contrastive prompting. Following the contrastive prompting method (Yao, 2025), we instruct Qwen3-8B to generate an easier and a more difficult rewrite at the same time and choose the difficult one as the final rewrite question. (3) SFT-only model and RL model without difficulty reward. We generate rewritten questions using the SFT-only model (without RL) and an RL model trained without the difficulty reward $R_{\text{diff}}$ as part of our ablation study. As shown in Table 2, our method significantly outperforms the rule-based and contrastive prompting on both metrics, demonstrating that RIDE rewrites achieve higher quality. Ablation study also shows the effect of our RL stage and difficulty reward.
Compared Method Dataset Average Consistent
Baseline RIDE Baseline RIDE
Rule-based AMC-23 40.83 58.75 28.33 48.33
AIME-24 36.67 63.33 28.89 55.56
Contrastive Prompting AMC-23 49.17 50.83 40.00 40.83
AIME-24 32.78 67.22 26.67 61.11
SFT-only AMC-23 38.75 60.00 33.33 53.33
AIME-24 36.67 62.22 26.67 52.22
w/o $R_{\text{diff}}$ AMC-23 43.75 54.17 35.83 47.50
AIME-24 27.78 72.22 21.11 65.56
Table 2: Win rate comparison of RIDE and other methods.
Case Study and Rewrite Strategy Analysis. Figure 4(a) illustrates the rewritten question. The original question is a classic coupon collector question with replacement. The rewritten question (1) clarifies the “pick 10 marbles at a time with replacement” procedure without changing the core idea; (2) raises the target success probability, increasing the required trials; and (3) requires using the Poisson approximation and the union bound, increasing restrictions. To systematically analyze these rewrites, we distill six core rewrite strategies from the questions in the RIDE-AIME and RIDE-AMC datasets: (1) Wording and Paraphrasing Modification, (2) Distracting Info, (3) Numerical Substitution, (4) Add Extra Steps, (5) Change Constraints, and (6) Change Target. (Detailed descriptions are provided in the Appendix A.4). Our analysis (Figure 4(b)) reveals that these strategies are frequently combined. Over 60% of the rewritten questions are composed of two or more strategies. To further investigate the impact on reasoning depth, we categorize strategies 11-33 as Surface Strategies, which have a minor impact on reasoning depth, and strategies 44-66 as Deep Strategies, which significantly affect the required reasoning. Statistics show that rewrites using a mix of both Surface and Deep strategies are the most common (44.3%). The accuracy of Qwen2.5-Math-72B and Llama4-Scout on questions corresponding to each strategy shows that performance varies across different strategies. In summary, our method can generate combinations of one or more rewrite strategies without explicitly defining specific rules. These combinations effectively provide a difficulty perturbation for the original question.
(a) Case study for our method. The original question is sampled from DeepMath dataset, and rewritten question is the corresponding rewritten version in RIDE-DeepMath.
(b) Rewrite strategy analysis.
Figure 4: Case study and analysis. (a) The original and rewritten questions from DeepMath and RIDE-DeepMath. (b) The analysis of rewrite strategies.
The Relation Between Difficulty and Robustness. To validate whether the performance drop stems from an increase in human-perceived difficulty or from the model’s inherent vulnerability, we conduct an analysis using the surface-level rewriting dataset mentioned above, which is assumed to keep human cognitive difficulty invariant. We record the transitions between “Correct on Original, Incorrect on Rewritten” and the inverse for each paired sample and apply McNemar’s test. The resulting p-value of $2.3847\times 10^{-11}$, being lower than $0.05$, reveals a significant asymmetry. This demonstrates that the performance drop is driven by the model’s non-robustness rather than shifts in difficulty. Additionally, as detailed in the Appendix B.1, we categorize questions into three difficulty levels and the consistent degradation across all difficulty levels (e.g., $11.43\%$ on ”Easy” questions and $12.86\%$ on ”Medium” questions) suggests this drop is independent of reasoning depth. From a theoretical perspective, the IRT-based difficulty is statistically estimated from the model’s response patterns, which is highly correlated with empirical accuracy. This difficulty focuses on the probability of a correct response rather than directly reflecting human cognitive complexity. Consequently, our “difficulty evolution” essentially transforms questions into formats that are challenging for the model to solve, thereby serving as a robustness test.
Data Augmentation For Training Data. Our method also serves as an effective data augmentation strategy for mathematical reasoning. Using RIDE-8B, we applied difficulty-evolution rewriting to the DeepMath dataset and filtered incorrect samples with DeepSeek-V3.1, yielding RIDE-DeepMath with over 10K examples. Without SFT, we trained Qwen3-0.6B, Qwen3-1.7B and Gemma3-4B-It via reinforcement learning with GSPO, producing RIDE-Qwen-0.6B-Thinking, RIDE-Qwen-1.7B and RIDE-Gemma-4B. Results (Table 3) show that models trained with reinforcement learning on our data achieve improvements over their base models. In particular, RIDE-Qwen-0.6B-Thinking and RIDE-Gemma3-4B have surpassed GPT-4o and Qwen2.5-Math-72B, and are close to Qwen3-4B in the non-thinking mode.
Model Params AIME-25
GPT-4o - 13.33
Qwen3-0.6B 0.6B 3.33
Qwen3-0.6B-Thinking 0.6B 10.00
RIDE-Qwen-0.6B-Thinking 0.6B 20.00
Qwen3-1.7B 1.7B 6.67
RIDE-Qwen-1.7B 1.7B 13.33
Qwen3-4B 4B 23.33
Gemma3-4B-It 4B 13.33
RIDE-Gemma3-4B 4B 20.00
Qwen2.5-Math-7B 7B 6.67
Qwen2.5-Math-72B 72B 13.33
Qwen3-30B-A3B 30B 23.33
Table 3: Pass@1 performance comparison on AIME-25 benchmark. Models trained with reinforcement learning on our augmented rewritten data achieve higher performance than the corresponding base models.
Effect Analysis. We additionally train RIDE-4B based on Qwen3-4B and keep SFT checkpoints (RIDE-4B-SFT, RIDE-8B-SFT) to disentangle SFT from RL effects. As shown in Table 4, Difficulty score is defined as the predicted probability from our difficulty ranker that the rewritten question is more difficult than the original. The experiments demonstrate that the RL-tuned RIDE-8B achieves the highest average Difficulty score, thereby validating the effect of our RL strategy. We also conduct ablation studies on the RIDE-4B model. We introduce two variations: (1) replacing the verifier for the correctness reward $R_{\text{cor}}$ with DeepSeek-V3.2 (RIDE-4B-DeepSeek), and (2) removing $R_{\text{cor}}$. We employ GPT-5 to evaluate the correctness of the rewritten questions generated by these models. The results indicate that RIDE-4B (GPT-5-mini verifier) achieves the highest correctness. RIDE-4B-DeepSeek model follows, as DeepSeek-V3.2 tends to output more false negatives (see Appendix B.4), whereas removing the correctness reward leads to decline in correctness. The DeepSeek verifier also improves correctness shows that our method is not heavily dependent on closed-source models, as open-source models are also effective. This confirms the validity and necessity of our correctness reward.
Besides, We compare different IRT models for question difficulty estimation. The 2-PL model adds a discrimination parameter to the 1-PL, while the 3-PL further introduces a guessing parameter. We randomly mask 20% of the response matrix for model fitting, and evaluate performance using AUC-ROC. We ultimately adapt the 1-PL model with VI for parameter estimation.
Part A: Effect analysis of RL and correctness reward
Model Difficulty Correctness
Qwen3-8B 74.61 40.00
RIDE-8B-SFT 78.78 40.00
RIDE-8B 80.18 42.50
RIDE-4B - 45.00
RIDE-4B-DeepSeek - 37.50
RIDE-4B w/o $R_{\text{cor}}$ - 32.50
Part B: Comparison of different IRT models
Model Method AUC-ROC
1-PL VI 91.29
1-PL MCMC 89.81
2-PL VI 86.56
3-PL VI 91.01
Table 4: Effect analysis. Part A shows the ablation study on RL and correctness reward $R_{\text{cor}}$, while Part B compares different IRT models and parameter estimation methods.
7 Conclusion
In this paper, we propose a question rewriting framework based on difficulty evolution. The method estimates question difficulty using IRT, then trains a difficulty ranker to provide a difficulty reward, guiding reinforcement learning for the rewriting model. Experiments show that our method causes performance drops in $26$ LLMs, validating its effectiveness and robustness in adversarial perturbation. Our method outperforms traditional rule-based perturbations and can be applied for training data augmentation. Further analyses confirm its effectiveness, offering a new perspective on robust evaluation from the view of “question difficulty”.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Ethics Statement
All data used in this work are from publicly available open source datasets under their respective licenses. For IRT modeling, we rely exclusively on responses generated by LLMs, and no real student data or personally identifiable information is used. Our method is designed only to rewrite mathematical questions and does not attempt to infer or profile individuals.
Reproducibility Statement
All our source code will be submitted in the Supplementary Materials, along with an explanation to facilitate reviewers in reproducing our work. The hyperparameter setup is introduced in Appendix A.2. We will also open source our mathematical question rewriting models RIDE-4B and RIDE-8B, our perturbation benchmarks RIDE-AMC and RIDE-AIME, and our augmented training data RIDE-DeepMath.
Appendix A Experimental Setup
A.1 Implementation Details
We use the API interface provided by OpenAI to call models and obtain the response matrix of LLMs (for the GLM series, we use the ZHIPU interface). The size of the response matrix is 3535×$2000$. We perform IRT parameter estimation through py-irt (Lalor and Rodriguez, 2023). The statistics of each student model’s performance is shown in Figure 5, along with the accuracy and estimated ability $\theta$ are shown in Table 5. The results indicate that within the same model series (e.g., Qwen2.5 or Qwen3), model parameter size is positively correlated with capability. However, this relationship is not numerically linear. Additionally, frontier reasoning models generally exhibit higher capability scores.
Model Parameter Pass@1 (%) $\theta$
Claude-3.5-haku - 46.75 -1.132
Claude-3.7-sonnet - 76.85 0.710
Deepseek-R1-Distill-Llama-8B 8B 74.65 0.552
Deepseek-R1-Distill-Qwen-1.5B 1.5B 71.65 0.324
Deepseek-R1-distill-Qwen-14B 14B 88.75 1.822
Deepseek-v3-0328 671B 85.85 1.534
Gemini-2.0-flash - 82.65 1.203
Gemini-2.5-flash - 93.15 2.406
Gemma3-12B-it 12B 64.80 -0.115
Gemma3-4B-it 4B 49.80 -0.970
GLM-4-9B-chat 9B 30.50 -2.038
GLM-4-plus - 60.25 -0.347
GLM-Z1-airx - 93.10 2.417
GLM-Z1-flash - 91.80 2.222
GPT-3.5-turbo - 29.85 -2.079
GPT-4o - 64.15 -0.135
Grok-2 - 62.35 -0.213
Grok-3 - 81.25 1.064
Kimi-K2-Instruct 1TB 93.05 2.462
Llama-3.1-70B 70B 53.10 -0.770
Llama-3.1-8B 8B 28.20 -2.168
Llama-4-scout 109B 70.90 0.284
Mistral-7B-Instruct-V0.2 7B 13.20 -3.427
Mistral-Small-3.1-24B-Instruct 24B 50.05 -0.947
Qwen-math-72B 72B 75.05 0.555
Qwen-math-7B 7B 63.60 -0.143
Qwen2.5-0.5B 0.5B 19.25 -2.799
Qwen2.5-14B 14B 64.90 -0.106
Qwen2.5-72B 72B 71.90 0.332
Qwen2.5-7B 7B 56.20 -0.541
Qwen3-0.6B 0.6B 26.95 -2.263
Qwen3-14B 14B 77.05 0.735
Qwen3-32B 32B 79.00 0.883
Qwen3-235B-A22B 235B 83.95 1.262
QwQ-32B 32B 95.00 2.813
Table 5: Statistics of each student model’s performance and estimated ability.
Figure 5: Statistics of each student model’s performance. Difficulty Level is given by DeepMath Dataset.
In the SFT stage, we use the LLaMA-Factory (Zheng et al., 2024) framework for training, with a dataset of $2,000$ rewritten mathematical questions distilled from GPT-5-mini. In the reinforcement learning stage, we use the TRL (von Werra et al., 2020) framework for training, with $3,000$ questions extracted from the DeepMath dataset (He et al., 2025).
A.2 Hyperparameter Setup
The hyperparameters used in our work are summarized in Table 6. For the IRT parameter estimation and the training of the difficulty ranker, we perform a greedy search to select the optimal group of hyperparameters based on evaluation metrics. For the SFT and reinforcement learning stages, the hyperparameters are chosen by balancing model performance with the available computational resources.
Hyperparameter Description Value
IRT & Difficulty Ranker
$M$ The number of student models 35
NN The number of mathematical questions 2,000
$S$ The number of virtual students by VAE 200
$T$ The number of virtual students by Sampling 200
$\beta$-VAE Weight of KLD loss for VAE 0.5
Latent dimension Dimension of latent space 32
Priors Prior distribution of IRT parameters vague
Random Seed Random seed for partitioning missing values in the response matrix 42
Batch size Number of samples per batch 64
Embedding dimension Dimension of input embeddings 3,072
Hidden dimension Hidden layer size of MLP (512,256)
Dropout rate Dropout probability for hidden layers 0.3
Activation Non-linear activation function ReLU
Epochs (CV) Training epochs for cross-validation 8 (5-CV)
Supervised Fine-Tuning
Device Device used for SFT 1 * NVIDIA A800 80GB PCIe
Finetuning type Finetuning strategy full
Template Prompt/format template of LLaMA-Factory qwen3
Cutoff length Max sequence length (tokens) 8192
Per-device train batch size Batch size per device 2
Gradient accumulation steps Grad accumulation steps 4
Learning rate Initial learning rate 1e-5
Epochs Number of training epochs 3.0
Warmup ratio Warmup ratio 0.1
Bf16 bfloat16 training true
Deepspeed (Rasley et al., 2020) Deepspeed Stage 3
Reinforcement Learning
Device Device used for RL 4 * NVIDIA A800 80GB PCIe
Finetuning type Finetuning strategy LoRA
Method RL algorithm GSPO
Per-device batch size Prompts per device per step 1
Gradient accumulation steps Steps to accumulate gradients 4
Group size $K$ Generations per prompt 4
Max prompt length Max tokens of prompt 512
Max new tokens Max tokens to generate per sample 4096
Temperature Sampling temperature 0.7
Top-p Nucleus sampling p 0.95
Max steps Maximum training steps 500
KL coef KL penalty coefficient 0.0
LoRA rank Rank 64
LoRA $\alpha$ Scaling factor 64
LoRA dropout Dropout probability 0.05
Deepspeed Deepspeed Stage 2
$\alpha$ Weight of difficulty reward 0.5
$\beta$ Weight of correctness reward 0.5
$\gamma$ Weight of keyword and length reward 0.3
Table 6: Main hyperparameter setup.
A.3 Prompt Strategy
(a) Prompt for mathematical question answering.
(b) Prompt for rewriting mathematical questions.
(c) Prompt for judging which rewritten question is better.
Figure 6: Overview of prompts.
For constructing the response matrix and evaluating model performance on both the original benchmark and the perturbed benchmark, we use the same zero-shot prompt to instruct models to answer mathematical questions (Figure 6(a)), requiring only that answers follow a specific format for verification. In practice, we find that even without explicitly telling the model to “think step by step”, all evaluated models still produce chain-of-thought outputs.
Figure 6(b) shows the prompt we use to guide the model in rewriting mathematical questions during the SFT and reinforcement learning stages. We do not impose a single, detailed rule. Instead, we briefly outline several ways to increase difficulty and allow the model to improvise, making the trajectory of difficulty evolution unpredictable.
Figure 6(c) presents the prompt we use to instruct GPT-5 to choose the one that better meets the requirements from two rewritten questions, enabling us to compute the win rate of our method against the rule-based approach. We specify in detail that the rewritten question must address covering question completeness, clarity, conceptual consistency, structural perturbations, and the avoidance of reasoning shortcuts.
A.4 Rewrite Strategy
We distill six core rewrite strategies from the rewritten questions: (1) Wording and Paraphrasing Modification, (2) Distracting Info, (3) Numerical Substitution, (4) Add Extra Steps, (5) Change Constraints, and (6) Change Target. The detailed descriptions are shown in Table 7. To determine the strategies involved in each question, we employed a model-based annotation process. We fed each question to both GPT-5 and DeepSeek-V3.2 three times each, for a total of six annotations. A specific strategy was only recorded as being present for a given question if it was selected by a majority (i.e., at least three out of six) of the model runs.
Strategy Definition
Wording Rephrases the question using different vocabulary or structures without altering math logic.
Distracting Adds irrelevant information, definitions, or values to the question description.
Numerical Replaces key numerical values in the question with new values.
Extra Steps Introduces new events or steps that must be incorporated into the calculation.
Constraints Adds stricter limitations or alters core rules to change the solution’s domain.
Target Changes the final expression to a more complex or generalized version.
Table 7: Definitions of the Rewrite Strategies.
Appendix B Effect Analysis
B.1 Orthogonality between question difficulty and robustness verification
The verification of model robustness and the evolution of question difficulty can coexist in our method. We define the robustness of LLMs in mathematical reasoning as follows: for a question $x$ and the model’s response correctness $y\in\{0,1\}$, given a semantic-preserving transformation family $\mathcal{T}$, if $f(x)=y$, then for the vast majority of $t\in\mathcal{T}$, $f(t(x))=y$ holds. Figure 8 shows that the semantic similarity between our rewritten and original questions consistently exceeds 80%, confirming that performance degradation does not stem from semantic disparity. Building on this, we decouple question difficulty from robustness verification. We identify ”surface-level modifications” that preserve reasoning depth. An example is shown in Figure 7.
Figure 7: A case study of the rewritten questions with surface-only strategies.
In this example, the evolution of the question is manifested in the formalized mathematical conditions and changes in phrasing, which do not impact the depth of reasoning required, making this a case for verifying model robustness. For example, Qwen2.5-72B answered the original question correctly but failed on the rewritten question. We collected similar surface-only rewrites and categorized them into three difficulty levels—easy, medium, and hard—based on the accuracy rates of $10$ LLMs on the original questions. The questions rewritten using the surface-only strategy maintain a reasoning depth consistent with the original questions, leaving the surface form as the sole variable. If the model’s performance degradation were solely dependent on difficulty, the magnitude of the performance drop on the rewritten questions should vary significantly across the three difficulty levels. Conversely, if question difficulty and robustness verification can be decoupled, the variations in performance decline across the three difficulty levels should be insignificant. We evaluated the performance degradation between the original and rewritten questions across different difficulty levels (Table 8).
Level Original Accuracy Rewritten Accuracy $\Delta$
Easy 80.00 68.57 11.43
Medium 68.57 55.71 12.86
Hard 62.50 43.75 18.75
Table 8: Performance drop comparison across different difficulty levels.
The results indicate that surface-level rewriting induces performance degradation even on easy questions, thereby highlighting the necessity of robustness verification. Moreover, the performance drop across difficulty levels is not significant, with a particularly negligible variation in the magnitude of change between easy and medium questions. We additionally recorded the counts of ”Correct on Original, Incorrect on Rewritten” and ”Incorrect on Original, Correct on Rewritten” for each paired sample in the surface-level rewriting dataset and calculated the p-value using McNemar’s test. The resulting p-value of $2.3847\times 10^{-11}$ is substantially lower than 0.05, revealing a significant asymmetry. This demonstrates that the performance decline originates from the model’s non-robustness rather than changes in difficulty. In summary, through case study and statistical analysis, we decoupled question difficulty from robustness verification and found that the two are orthogonal, allowing them to coexist within our method.
B.2 Effect of IRT models and difficulty ranker.
As a limited number of student models is not enough to estimate IRT parameters accurately, while employing a large number of LLMs as students is computationally costly, we augment the student–response data to simulate additional student behaviors. We adopt two augmentation strategies:
  • Variational Autoencoder (VAE) method. We mitigate the scarcity of student responses by training a VAE (Kingma and Welling, 2022) on the observed response matrix $\mathbf{Y}$ to learn a low-dimensional manifold of student behavior. During generation, we draw a latent code $\mathbf{z}$ from the Gaussian prior and decode it into per-question correctness probabilities $\pi(\mathbf{z})$, from which Bernoulli draws yield a synthetic student-question response vector with conditional independence across items given $\mathbf{z}$:
    $\mathbf{r}_{\text{VAE},j}\mid\mathbf{z}\sim\mathrm{Bernoulli}\!\big(\pi_{j}(\mathbf{z})\big),\qquad j=1,\dots,N$
  • Sampling method. We simulate the responses of students by drawing from the empirical distribution observed in $\mathbf{Y}$. For each question $j$, we compute the empirical correctness rate $C$ over the observed responses and use it as the Bernoulli success probability:
    $\mathbf{r}_{\text{Sampling},j}\sim\mathrm{Bernoulli}\!\Big(C_{j}\Big),\qquad j=1,\dots,N$
    Each sampled $\mathbf{r}_{\text{VAE}}$ and $\mathbf{r}_{\text{Sampling}}$ is then appended as a new student row to $\mathbf{Y}$.
These methods let us synthesize student–item response vectors that respect observed item statistics and student tendencies. We concatenate the synthetic responses $\mathbf{r}_{\text{VAE}}$ and $\mathbf{r}_{\text{Sampling}}$ with the originals to form an augmented matrix $\tilde{\mathbf{Y}}\in\{0,1\}^{M+S+T,N}$:
$\tilde{\mathbf{Y}}$ $=\begin{bmatrix}\mathbf{Y}\\ \mathbf{R}_{\mathrm{VAE}}\\ \mathbf{R}_{\mathrm{Sampling}}\end{bmatrix}$, (1)
$\mathbf{R}_{\mathrm{VAE}}$ $=[\mathbf{r}_{\mathrm{VAE}}^{1},\dots,\mathbf{r}_{\mathrm{VAE}}^{S}]^{\intercal},$
$\mathbf{R}_{\mathrm{Sampling}}$ $=[\mathbf{r}_{\mathrm{Sampling}}^{1},\dots,\mathbf{r}_{\mathrm{Sampling}}^{T}]^{\intercal}$
$S$ and $T$ denote the number of synthetic students by VAE and sampling, respectively. Then the modeling ability of IRT calibration for question difficulties can be improved. We perform IRT parameter estimation based on the response matrix $\tilde{\mathbf{Y}}$.
As described in Section 6, we conducted additional experiments to verify the effectiveness of the choice of the IRT model and data augmentation methods. The results are shown in Table 9. We consider two priors provided by py-irt: a vague prior that exerts minimal influence so the likelihood largely drives estimation, and a hierarchical prior that treats parameters as draws from a shared group-level distribution with hyperpriors, enabling partial pooling to stabilize estimates and shrink extremes, especially with sparse data. For evaluation, in addition to AUC-ROC for discrimination, we also report the Brier Score, the mean squared difference between predicted probabilities and observed outcomes (lower is better), which captures overall probabilistic accuracy and calibration and thus complements AUC-ROC’s ranking focus.
[…]
VAE Sampling
1-PL VI - - vague 89.81 11.97
1-PL VI - - hierarchical 90.03 11.89
1-PL VI - vague 91.01 11.54
1-PL VI - hierarchical 90.95 11.61
1-PL VI - vague 91.06 11.25
1-PL VI - hierarchical 91.03 11.29
1-PL VI vague 91.29 11.29
1-PL VI hierarchical 91.27 11.32
1-PL MCMC - - vague 89.81 11.95
1-PL MCMC - - hierarchical 90.08 11.85
2-PL VI - - vague 85.18 15.87
2-PL VI - - hierarchical 90.41 11.72
2-PL VI - vague 86.19 15.50
2-PL VI - hierarchical 90.89 11.64
2-PL VI - vague 86.24 15.73
2-PL VI - hierarchical 89.62 12.53
2-PL VI vague 86.56 15.52
2-PL VI hierarchical 89.81 12.27
2-PL MCMC - - vague 84.56 15.72
2-PL MCMC - - hierarchical 88.06 13.34
3-PL VI - - vague 89.52 12.21
3-PL VI - - hierarchical 89.52 12.21
3-PL VI - vague 90.76 11.73
3-PL VI - hierarchical 90.76 11.73
3-PL VI - vague 90.05 12.50
3-PL VI - hierarchical 90.05 12.50
3-PL VI vague 91.01 12.13
3-PL VI hierarchical 91.01 12.13
Table 9: Comparison of IRT models, data augmentation and prior method.
Besides, we explore regression, classification, and ranking methods to fit IRT-estimated question difficulties from textual features. Regression task is evaluated by RMSE, consistently yields values above $1$ across models and embeddings, indicating large bias. Classification treats questions with difficulty below 0 as “easy” and above 0 as “hard,” but F1 scores remain below 6060% due to class imbalance. Finally, we adopt a pairwise ranking approach with AUC-ROC as the metric, achieving around 7272%, with little variation across embedding models.
Method Task Embedding Metrics
LR Regression Qwen3-8B 1.0375
LR Regression BGE-M3 (Chen et al., 2024) 1.1380
SVR Regression Qwen3-8B 1.0108
MLP Regression Qwen3-8B 1.0931
Longformer (Beltagy et al., 2020) Regression Qwen3-8B 1.0746
SVM Classification Qwen3-8B 58.34
MLP Classification Qwen3-8B 53.60
MLP Ranking Qwen3-8B 71.95
MLP Ranking text-embedding-3-large 71.98
Table 10: Comparison of regression, classification and ranking tasks on different models.
B.3 Statistics of RIDE rewritten model.
Starting from Qwen3-1.7B and Qwen3-4B, we trained RIDE-1.7B and RIDE-4B, and preserved SFT checkpoints (RIDE-1.7B-SFT, RIDE-4B-SFT, RIDE-8B-SFT) to separate SFT from RL effects. The questions rewritten by each model are evaluated in terms of average generation time, average token count, correctness, and semantic similarity. Generation time and token count use the minimum of three runs, correctness is judged by GPT-5 and similarity is measured via text-embedding-3-large. The results are shown in Figure 8. Compared to Qwen3-8B, our SFT and reinforcement learning models produce longer outputs and slower inference, showing that rewrites expand questions, solutions, and answers. Semantic similarity drops slightly, indicating greater diversity while preserving affinity with the original questions. Reinforcement learning improves correctness over both the original and SFT-only models. RIDE-8B and RIDE-4B demonstrate much higher correctness than RIDE-1.7B, indicating that parameter size influences the correctness of generated questions. However, this difference is not significant between the 4B and 8B models, suggesting a limited effect of scale within this range.
Figure 8: Statistics of average token count, average generation time, correctness and semantic similarity across different models.
B.4 Effect of LLM verifier
We use the proprietary GPT-5 series models to provide correctness rewards and as filtering models. To verify that our method can also use low-cost, open-source models, we employ GPT-5-mini and Deepseek-V3.2 (non-thinking mode) as verifiers. We provide $2,000$ math questions (without answers), along with the solutions and answers generated by two test models (Qwen2.5-Math-7B and Gemma3-12B), as input to these verifiers to judge the correctness of the test models’ responses. The results, shown in Table 11, indicate that the verification accuracy of the closed-source model GPT-5-mini reach over 85%. In contrast, the accuracy of DeepSeek-V3.2 was relatively lower, but still over 65%. However, both verifiers achieve a precision of around 90%. This indicates that both models can precisely assign low reward signals to incorrect solutions and are effective for filtering erroneous data. The relatively low recall, however, implies that the verifiers tend to be conservative, flagging an otherwise correct solution and answer as incorrect. This may be because the model’s solution contains minor errors even when the final answer is correct. This situation is acceptable, given that rule-based or model-based rewards in Reinforcement Learning with Verified Rewards (RLVR) also exhibit a high number of false negatives (Xu et al., 2025a). While this situation affects training efficiency, it does not directly lead to training failure, as our verifier assigns low reward signals to incorrect rewrites. Therefore, our method can also use open-source models as verifiers and is not merely a distillation from proprietary models.
Response Source Verifier Precision Recall F1 Accuracy
Qwen2.5-Math-7B GPT-5-mini 97.96 79.17 87.57 85.70
Deepseek-V3.2 89.48 62.19 73.38 71.30
Gemma3-12B GPT-5-mini 98.38 79.71 88.06 86.00
Deepseek-V3.2 93.84 54.09 68.62 67.95
Table 11: Verification performance of GPT-5-mini and Deepseek-V3.2 verifiers on solutions generated by different response models.
Appendix C Item Response Theory Analysis
C.1 Human Difficulty & LLM Difficulty
In human cognition, the difficulty of a math question is often determined by its required reasoning depth. Consequently, researchers have established the AoPS (Art of Problem Solving) standard to measure the difficulty of math competition questions from a human perspective, and methods have emerged that use LLMs in conjunction with the AoPS standard to automatically assign difficulty ratings to math questions  (Gao et al., 2025). The DeepMath dataset, which we utilize in the IRT modeling phase, also applies this standard to provide difficulty levels for its questions. However, we discover a discrepancy between this human-centric standard and actual LLM performance. After grouping the responses of our $35$ models on $2,000$ questions by their given AoPS difficulty levels, we calculate the mean empirical accuracy for each level. The Spearman and Kendall correlation coefficients between the AoPS difficulty and the mean empirical accuracy are found to be -0.5152 and -0.5111, respectively, indicating that while the expected negative correlation exists (i.e., higher human-perceived difficulty corresponds to lower model accuracy), the relationship is not sufficiently strong to consider them equivalent. Therefore, we conclude that a deviation exists between the human cognitive difficulty defined by the AoPS standard and the empirical error rate demonstrated by LLMs. Furthermore, treating lower empirical accuracy directly as higher ”LLM difficulty” introduces a fundamental flaw: empirical accuracy assigns equal weight to the responses of all tested LLMs. We therefore introduce IRT to define difficulty, which allows us to derive a statistically robust difficulty value by incorporating the ability of all participating LLMs into the calculation.
Additionally, we compare the correlation of both the empirical error rate (ungrouped) and IRT difficulty against the human-perceived difficulty, again using Spearman and Kendall coefficients, as shown in Table 12. The results show that both empirical error rate and our IRT difficulty (fitted on 35 model responses) have a similar correlation with the given human difficulty. Our work is primarily focused on this ”LLM difficulty” rather than human-perceived difficulty.
vs Human-Rated Difficulty Spearman Kendall
Empirical Error Rate 0.3362 0.2470
IRT Difficulty 0.3418 0.2468
IRT Difficulty w/o augmentation 0.3364 0.2429
IRT Difficulty (Qwen-only) 0.2839 0.2031
Table 12: Correlation analysis of empirical error rate, IRT difficulty and human-rated difficulty. Spearman & Kendall measure the strength and direction of the monotonic relationship (or rank-order correlation) between two variables.
We also conduct ablation studies by fitting IRT models on the original response matrix without augmentation, and on a matrix using only the Qwen-series models. The results show that the difficulty values derived solely from the Qwen series ($11$ test-takers) correlate poorly with both empirical accuracy and human difficulty, as the limited number of subjects introduces significant bias into the IRT fit. Besides, the original response matrix without augmentation is also less correlated due to data sparsity. This finding validates the effectiveness of our VAE and Sampling data augmentation methods. By learning from the original response matrix and empirical accuracy to generate a large number of distributionally-consistent virtual responses to expand the matrix, our approach enhances the robustness of the IRT fitting process. This prevents parameter estimation bias caused by an insufficient number of subjects.
C.2 Case for IRT Difficulty
As previously discussed, although both empirical error rate and IRT difficulty can be regarded as values for LLM Difficulty, empirical error rate’s drawback is its assumption of equal weighting for all LLMs, leading to a lack of item discrimination. In Figure 9, we present a case study of two questions. Although both questions have the identical empirical accuracy (37.14%), their IRT difficulties differ significantly: Question 1 has an IRT difficulty of 0.599, whereas Question 2 has an IRT difficulty of 0.778. This discrepancy arises because the profile of models that answer each question correctly is different. Specifically, four high-ability models including Deepseek-R1-Distill-Qwen-1.5B, Gemini-2.5-Flash, GLM-Z1-AIRX, and Qwen3-235B-A22B (see ability values in Table 5) answer Question 1 correctly but fail to answer Question 2. Therefore, even with the same empirical accuracy, Question 1 is more likely to be solved by high-ability LLMs, thus receiving a lower IRT difficulty score. This case validates the advantage of using IRT difficulty over simple empirical error rate, as it successfully distinguishes item difficulty based on the ability of the solvers, not just the raw pass count.
Figure 9: A case study of two questions with identical empirical accuracy and different IRT difficulties.
C.3 Distribution Shift
To validate the stability of our IRT parameter estimations, we collect additional responses for the same $2,000$ items using five extra models: Qwen2-72B, Mistral-large, GPT-oss-20B, Qwen3-1.7B, and GPT-4.1-nano. Among these, Qwen2-72B and Qwen3-1.7B are classified as ”lower-performing LLMs” as their accuracies fell below the $67.98\%$ average accuracy of our initial $35$ LLMs. We then conduct two experiments: first, we added only these weaker LLMs to the original $35$-LLM response matrix; second, we added all $5$ new models to the original 35-LLM matrix. We then fit IRT parameters for both augmented matrices and calculate the Spearman and Kendall correlation coefficients against the original IRT difficulty parameters derived from the initial $35$ LLMs. The results (shown in Table 13) indicate that in both scenarios—whether adding only the weaker models or all five additional models—the changes in the Spearman and Kendall correlation coefficients are minimal. This demonstrates the stability of our method, confirming that our IRT parameter estimations are robust and not significantly perturbed by the inclusion of a certain amount of new response data.
vs $35$-LLMs IRT Difficulty Spearman Kendall
$35$-LLMs + lower-performing $2$-LLMs 1.00 0.9968
$35$-LLMs + $5$-LLMs 1.00 0.9954
Table 13: Correlation analysis between the original $35$-LLM fitted IRT difficulty and those fitted with 22 additional lower-performing LLMs or all 55 additional LLMs.